# Zero-shot Hierarchical Plant Segmentation via Foundation Segmentation Models and Text-to-image Attention

### v1

@@ 4.5 Results
The domain gap of different shooting environments, such as different equipment or locations, also greatly influences the segmentation results, as the environment setups would significantly change the image quality. The evaluation and Visualization of the test on SB20, a dataset of different shooting environments (i.e., UAV and ground vehicle, different farms) for the same plant species as PhenoBench, are shown in Tab. 4 and Fig. 7, respectively. Since SB20 only provides the ground-truth data for plant individual segmentation, we do not evaluate leaf instance segmentation accuracies. So LO-SAM is not involved in this experiment. Compared with the zero-shot baseline, i.e., Grounded SAM, our method yields higher scores in most of the metrics. HAPT scores a relatively high value, but not as high as Grounded SAM and ours, and Weyler et al. almost always fail to segment crops. These results indicate that the current state-of-the-art methods relying on training datasets (i.e., HAPT and Weyler et al.) do not generalize to the different shooting environments, even with the same plant species.

### v2

@@ 4.5 Results
SB20 is a dataset with different shooting environments (i.e., UAV and ground vehicle, different farms) and the same plant species as PhenoBench. The accuracy and visualization of the test set on SB20 are shown in Tab. 4 and Fig. 7, respectively. Since SB20 only provides the ground-truth data for plant individual segmentation, we do not evaluate leaf instance segmentation accuracies, where LO-SAM is not involved in the experiment for this reason.
Compared with the zero-shot baseline, i.e., Grounded SAM, our method yields higher scores in most of the metrics. HAPT shows a relatively high accuracy, but not as high as Grounded SAM and ours, and Weyler et al. almost always fail to segment crops. These results indicate that the current state-of-the-art methods relying on training datasets (i.e., HAPT and Weyler et al.) do not generalize to the different shooting environments, even with the same plant species. The domain gap of different shooting environments greatly influences the segmentation results for the supervised method, which implies the impracticability of current state-of-the-art methods.
@@ Acknowledgments
Appendix A Technical Details
A.1 Cross-Attention and Weighted Least Square Line
In this section, we describe the text-to-image cross-attention that we use to obtain the center-like probability of each leaf mask and the weighted least square line that we use its edge points as key points of each leaf.
Text-to-Image Cross-Attention
Text-to-image cross-attention in one head is described as Eq. (5), where $\bm{F}_{I}\in\mathbb{R}^{N_{I}\times d}$ and $\bm{F}_{T}\in\mathbb{R}^{N_{T}\times d}$ denote the image feature and the text feature, respectively.
$\text{CrossAttention}(\bm{F}_{I},\bm{F}_{T})=\text{softmax}\left(\frac{\bm{F}_{I}\bm{F}_{T}^{\mathsf{T}}}{\sqrt{d}}\right).$~ (5)
In their official implementation  [^1] Grounding DINO, https://github.com/IDEA-Research/GroundingDINO, last accessed on July 10, 2025., the cross-attention module of Grounding DINO [16] consists of six cross-attention layers. We use the output of the final layer to obtain the feature map. Specifically, we extract four feature maps in resolution from $h/8\times w/8$ to $h/64\times w/64$ from each head, where $(h,w)$ denotes the size of the input image to the module. We resize to $h/8\times w/8$ and average them to obtain the feature map of the $i$-th leaf mask. We describe it as $f^{i}\in\mathbb{R}^{\frac{h}{8}\times\frac{w}{8}}_{+}$
Algorithm 1 Greedy DBSCAN
[…]
3:function GreedyDBSCAN($\hat{\mathcal{Q}}$, $\epsilon$, MinPts)
4:  $\mathcal{C}\leftarrow\varnothing$
5:  $\text{MinPts}\leftarrow\text{InitMinPts}$
6:  for (j=1;j<num_max_cluster;j+=1)(j=1;j<num\_max\_cluster;j+=1) do
7:   $\tilde{\mathcal{C}}\leftarrow\mathrm{DBSCAN}(\hat{\mathcal{Q}},~\epsilon,~\text{MinPts})$
8:   if $\tilde{\mathcal{C}}==NULL~\textbf{and}~\text{MinPts}==2$ then
9:     break
10:   end if
11:   $C^{j}\leftarrow\mathop{\arg\max}_{C\in\tilde{\mathcal{C}}}|C|$
12:   $\hat{\mathcal{Q}}\leftarrow\hat{\mathcal{Q}}-C^{j}$
13:   $\mathcal{C}\leftarrow\mathcal{C}\cup C^{j}$
14:   if $\text{MinPts}>2$ then
15:     $\text{MinPts}\leftarrow\text{MinPts}-1$
16:   end if
17:  end for
18:  return $\mathcal{C}$
19:end function
Weighted Least Squares Line
Now, we have the feature map of the $i$-th leaf mask, $f^{i}$. To determine the key points of the mask, we calculate the weighted least squares (WLS) line on the feature map and obtain key points as its edge points. Considering the value at the coordinate $(a,b)$ in the feature map $f^{i}[a,b]$ as the weight at the point $(a,b)$, an objective function of the WLS problem is described as Eq. (6), where $\bm{x}=[x_{1},x_{2}]^{\mathsf{T}}$ describes the WLS line.
[…]
Solving Eq. (6) with $\frac{\partial J(\bm{x})}{\partial x_{1}}=\frac{\partial J(\bm{x})}{\partial x_{2}}=0$, the WLS line $\bm{x}$ can be calculated as Eq. (7). For more detailed explanations, please refer to [5].
$\bm{x}=(\bm{A}^{\mathsf{T}}\bm{W}\bm{A})^{-1}\bm{A}^{\mathsf{T}}\bm{W}\bm{b}.$ (7)
A.2 Greedy Clustering
Algorithm 1 shows the pseudo-code of the greedy clustering for plant individual segmentation.
Figure 8: Comparison of ZeroPlantSeg and LeafOnlySAM. Since LeafOnlySAM does not involve plant segmentation, we only show the leaf instance segmentation results on the PhenoBench and Growliflower datasets. The left three columns show the results on the GrowliFlower dataset, and the right three columns show the results on the PhenoBench dataset.
Figure 9: Comparison of ZeroPlantSeg and Grounded SAM 2. The results are divided into two parts. The top rows show the comparison on the GrowliFlower dataset, and the bottom rows show the comparison on the PhenoBench dataset. Each picture is denoted by “Plant” or “Leaf” to indicate whether it is a leaf instance segmentation or a plant individual segmentation result.
Figure 10: Comparison of ZeroPlantSeg and Grounded SAM 2 on the SB20 dataset. Since SB20 only supplies annotations of individual plants, we only visualize the plant segmentation results.
Appendix B Additional Results
While our method quantitatively surpasses these existing methods, Figs. 8,  9, and 10 show the visual results for the recent zero-shot models, namely, Grounded SAM2 [22] and LeafOnlySAM [31], compared to ours. Similar to the trend shown in the quantitative evaluation, our method more accurately estimates leaf and plant individual instances than the state-of-the-art zero-shot segmentation models.